# Arrow-Debreu Meets Kyle: Price Discovery for Derivatives

### v26

@@ 9 Conclusion
By unifying the elements of Arrow and Debreu (1954) and Kyle (1985), we have developed a general model of informed trading across contingent claim markets. Our findings advance the basic theoretical understanding of how information regarding general aspects of underlying payoff distributions is incorporated into derivatives prices. By characterizing the endogenous interaction of information across derivative securities, our results provide explanations for observed empirical practices, bridge gaps with existing empirical literature, and propose new avenues for future empirical research.
@@ A.1 Model Assumptions
The pricing kernel follows. This proves the theorem.
@@
The verification of $\partial\!\kern 0.80002ptJ_{p}$ is immediate because the Gâteaux derivative of the functional
$W\,\stackrel{{\scriptstyle\partial\!\kern 0.56001ptJ_{p}}}{{\mapsto}}\,\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k}),\,\mathbb{R}^{\infty}\rightarrow\mathbb{R},$
is trivially identified with $\eta(\,\cdot\,|s_{k})$.
@@
For $c\colon S\rightarrow{\mathbb{R}}$, define the affine subspace of portfolios $\mathcal{V}_{c}(\widetilde{W})\!\equiv\!\{W\colon\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=c\}.$ The zero price impact subspace $\mathcal{V}_{0}(\widetilde{W})$ corresponds to $c=0$.
@@
Similarly, the transformation (22) replaces ${\bf L}^{2}$ and $d$ in the insider’s objective function $J(d|k;\widetilde{D})$ of (A.10) by ${\bf I}$ and $\hat{d}$, respectively. This verifies the insider’s problem (25) in the canonical game and proves the proposition.

### v27

@@ 9 Conclusion
By unifying the elements of Arrow and Debreu (1954) and Kyle (1985), we have developed a general model of informed trading across contingent claim markets. Our findings advance the theoretical understanding of how contingent claim markets jointly incorporate information. Our framework provides explanations for observed empirical practices, bridges the gap with empirical literature, and proposes new avenues for future research.
@@ A.1 Model Assumptions
This proves the theorem.
@@
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$ at $W$. (Like $J(\,\cdot\,|s_{k};\widetilde{W})$, its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ is understood to be conditional on $s_{k}$ and parameterized by $\widetilde{W}$.) We need to show this Gâteaux-derivative decomposes into
$\underbrace{\partial\!\kern 0.80002ptJ(\,\cdot\,;W)}_{\text{marginal utility}}=\underbrace{\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)}_{\text{marginal payoff}}-\underbrace{\big{(}\,\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(\,\cdot\,;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(\,\cdot\,;W)\,\big{)}}_{\text{marginal cost}}$
Because the Gâteaux derivative of the payoff functional
$W\,\mapsto\,\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k}),\,\mathbb{R}^{\infty}\rightarrow\mathbb{R},$
is trivially identified with $\eta(\,\cdot\,|s_{k})$, the characterization of the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ in Equation (15) is immediate.
@@
For $c\colon S\rightarrow{\mathbb{R}}$, define the affine subspace of portfolios $\mathcal{V}_{c}(\widetilde{W})\!\equiv\!\{W\colon\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=c\}.$ The zero price impact portfolios are those in the subspace $\mathcal{V}_{0}(\widetilde{W})$ corresponding to $c=0$.
@@
Similarly, the canonical transformation replaces ${\bf L}^{2}$ and $d$ in the insider’s objective function $J(d|k;\widetilde{D})$ of (A.10) by ${\bf I}$ and $\hat{d}$, respectively. This verifies the insider’s problem (25) in the canonical game and proves the proposition.